OpenStudy (anonymous):

@Mertsj Please help me with this! If $1000 is invested at 12% compounded monthly for 7 years the compounded amount is given by A=1000(1.01)^84. Given that log1.01=0.00432, find log A. (Note that 1000=10^3) Choices: a. 3.36288 b. 3.00432 c. 4.36288 d. 4.01 Thank you! All help is really appreciated =)

OpenStudy (mertsj):

\[\log A = \log 1000+\log1.01^{84}=3+84\log1.01=3+.362995398\]

OpenStudy (anonymous):

Thank you!

OpenStudy (mertsj):

yw